What is the wavelength of radui wave with a frequency of 1.5x10⁷Hz?



1
Expert's answer
2021-11-18T08:25:02-0500

Wavelength = λ = c/f; where f is the frequency in Hertz and c = 299792458 m/s


"Wavelength =\frac{ 299792458}{1.5×10^7}=19.97" m